A rectangle has a length of 11 inches less than 3 times its width. If the area of the rectangle is 592 square inches, find the length of the rectangle.
Found 3 solutions by fractalier, stanbon, It is costly:
Answer by fractalier(6550)

Perimeter P = 2L + 2W = 592 and we also have
L = 3W - 11
Substitute that into the first equation and get
2(3W - 11) + 2W = 592
6W - 22 + 2W = 592
8W - 22 = 592
8W = 614
W = 76.75
L = 3(76.75) - 11 = 219.25 inches

Answer by stanbon(75887)

A rectangle has a length of 11 inches less than 3 times its width. If the area of the rectangle is 592 square inches, find the length of the rectangle.
-----
width:: w
length:: 3w-11
-----
Equation:
Area = width*length
A = w(3w-11)
3w^2 - 11w = 592
3w^2 - 11w - 592 = 0
-----
width = 16; length = 3*16-11 = 48-11 = 37 inches
------------
Cheers,
Stan H.
----------

Answer by It is costly(175)